OpenStudy (anonymous):

Determine the factors of x2 − 10x + 24. (x + 6)(x + 4) (x − 3)(x − 8) (x + 3)(x + 8) (x − 6)(x − 4)

OpenStudy (goformit100):

Do you know how to find the factor of a quadratic expression ?

OpenStudy (anonymous):

nope

OpenStudy (goformit100):

first try to break 24 in such two halves so that you get 10 as answer

OpenStudy (anonymous):

i dont know how

OpenStudy (goformit100):

take out prime facto0rize 24 and say what r you getting ?

OpenStudy (anonymous):

i still dont know

OpenStudy (akashdeepdeb):

I have told you MANY times!! And you EVEN said You Understood me!!! :/

OpenStudy (anonymous):

okay but dont get mad

OpenStudy (anonymous):

okay but the problem is i dont understand it

OpenStudy (zzr0ck3r):

what are two numbers that add to 10 but multiply to 24?

OpenStudy (anonymous):

12 and -2

OpenStudy (zzr0ck3r):

that multiplies to -24

OpenStudy (anonymous):

yep

OpenStudy (zzr0ck3r):

we want it to multiply to 24 not -24

OpenStudy (zzr0ck3r):

two different numbers?

OpenStudy (anonymous):

12 and 2

OpenStudy (zzr0ck3r):

does that add to 10?

OpenStudy (zzr0ck3r):

try other numbers

OpenStudy (anonymous):

idk

OpenStudy (zzr0ck3r):

two numbers that multiply to get 24 how else can two numbers multiply to get 24 besides 2*12

OpenStudy (anonymous):

1 and 24

OpenStudy (zzr0ck3r):

what else

OpenStudy (zzr0ck3r):

1*2=24 2*12 = 24 ?*? = 24

OpenStudy (anonymous):

3 and 8 4 and 6

OpenStudy (zzr0ck3r):

which one of those add to 10?

OpenStudy (zzr0ck3r):

cmon I got to go soon

OpenStudy (anonymous):

4 an 6

OpenStudy (zzr0ck3r):

ok, now notice that we have -10 and not 10 as our middle term, but just as 4+6 = 10 and 4*6 = 24, we also have -6+-4 = -10 and -4*-6 = 24 so our numbers are -4,-6

OpenStudy (zzr0ck3r):

and since we dont have a coefficient on the first term, that is the answer

OpenStudy (anonymous):

oh okay thank you

OpenStudy (zzr0ck3r):

you should be able to list the combinations of factors to get numbers \[24=1*24=2*12=3*8=4*6=-4*-6=-3*-8=-2*-12=-1*-24=24\]

OpenStudy (zzr0ck3r):

\[=-1*-24=24\]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

its d right

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

you should be able to expand D) and know this for sure, if not...I would really suggest that you go back a few sections/chapter and catch up. Its only going to get harder, and it is 50000000 times harder if you dont understand the material that came before, as we are baring witness too.